The product of two consecutive numbers is 210
Answer by checkley77(12569)

X(X+1)=210
X^2+1X-210=0
(X-+15)(X-14)=0
X-14=0
X=14 ANSWER.
14+1=15 ANSWER.
PROOF:
14*15=210
210=210